# revtex4-2Repair the float

### v4

@@ C.1 Dynamics in Small Systems
Figure 12: Eigenvector statistics for Trotterised A2A-Ising unitary. Columns show increasing $j$ values from left-to-right, and the rows, from top-to-bottom, are for step sizes before the threshold ($\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$), chaotic ($\tau=0.5\leavevmode\nobreak\ (2\pi g^{-1})$), and stable revival (seen only in large system sizes) ($\tau=0.7\leavevmode\nobreak\ (2\pi g^{-1})$) regions. The coloured lines are RMT distributions as shown by the legend.
@@ C.2 Eigenvector Statistics
In the main text of the paper, we show eigenvector statistics histograms for $j=64$ and calculate $X^{2}_{\rm RMT}$ for the other system sizes. Here, to illustrate how system size affects the histograms themselves, we provide example histograms for several $j$ values and three step sizes, which are chosen from regular (before the threshold), quantum chaotic (after the threshold), and stable revival regions. Figure 12 shows the eigenvector statistics for different $j$ values on each column and different step sizes on each row: (top to bottom) before the threshold ($\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$), quantum chaotic ($\tau=0.5\leavevmode\nobreak\ (2\pi g^{-1})$), and stable revival (seen only in large system sizes) ($\tau=0.7\leavevmode\nobreak\ (2\pi g^{-1})$) regions, respectively. The smallest system size, in Fig. 12 (a), does not have enough components to produce meaningful statistics, and the visual comparisons in Figs 12 (b–c) do not provide very clear conclusions. The $X^{2}_{\rm RMT}$ goodness-of-fit test statistic, on the other hand, provides consistent conclusions in terms of successfully identifying the disagreement/agreement of eigenvector statistics with RMT distributions (shown in Fig. 6 (a,d)). For the large system sizes, the regular and chaotic regions are clearly distinguishable in the eigenvector statistics, shown respectively for (d) $j=64$ and (e) $j=256$ in Figs 12 (d–e). However, the stable regions ($\tau=4.5\leavevmode\nobreak\ (2\pi g^{-1})$) are visually indistinguishable from the quantum chaotic $\tau=3.5\leavevmode\nobreak\ (2\pi g^{-1})$ case, and, once again, $X^{2}_{\rm RMT}$ proves useful and identifies these weak islands (see main text).
@@ C.3 Stable Regions
The quantum-classical correspondence principle suggests the following explanation for the changes observed in the local observable in Fig. 13 (f) around $\tau=0.7\leavevmode\nobreak\ (2\pi g^{-1})$: Trotter errors are inevitably large for trajectories corresponding to chaotic dynamics that is ergodic within the entire classical phase space (as for a completely mixed state). Trajectories within the stable islands of classical phase space follow regular dynamics that is generally restricted to a more localised region, and which therefore exhibits a different overlap with the target dynamics, from that corresponding to an underlying classically chaotic trajectory. Hence the re-emergence of stable islands in the A2A-Ising’s classical phase space accompanies the revival of stability in the stable regions of the Trotterised quantum model, for initial states which overlap with the stable islands. But note again with caution that, as in the case of the Trotterised Heisenberg model, the observed revivals of stability do not typically correspond to any revival in the quality of simulation with an underlying target model.
@@ Appendix D Heisenberg Model
Figure 14: Eigenvector statistics for Trotterised Heisenberg unitary. Columns, left to right, show 3, 6, and 9 qubit chains, and rows, top-to-bottom, are for step sizes before the threshold ($\tau=2e\text{-}6\leavevmode\nobreak\ (2\pi g^{-1})$), quantum chaotic ($\tau=0.2\leavevmode\nobreak\ (2\pi g^{-1})$), and stable revival ($\tau\approx 0.5\leavevmode\nobreak\ (2\pi g^{-1})$) regions. 6 qubits and 9 qubits show clear deviations from the RMT distributions (c) before the threshold and (i) in the stable revival region, and it closely follows the COE distribution in (f) quantum chaotic region. 3 qubits cases in (a,d,g) do not have enough components to give good statistics. The coloured lines are RMT distributions as shown by the legend.
@@ D.1 Eigenvector Statistics
So far, we have only presented the processed $X^{2}_{\rm RMT}$ analysis for the Heisenberg Trotterisation eigenvector statistics. Here, Fig. 14 explicitly shows the full histogram data for 3, 6, and 9 qubits Heisenberg Trotterisations on each column and different step sizes on each row. The step sizes shown in the columns of Fig. 14 are, respectively from top-to-bottom, from before the threshold ($\tau=2e\text{-}6\leavevmode\nobreak\ (2\pi g^{-1})$), and from the quantum chaotic ($\tau=0.2\leavevmode\nobreak\ (2\pi g^{-1})$) and stable revival ($\tau\approx 0.5\leavevmode\nobreak\ (2\pi g^{-1})$) regions. (Note: Numerical diagonalisation becomes unstable at or near $\tau=0.5$ for our model parameters due to numerical imprecision, so we instead use $\tau=0.5000001$, which our analysis indicates is close enough to have “converged” to a limiting value for $\tau\rightarrow 0.5$, but far enough from $\tau=0.5$ for numerical error to be neglected.) Unlike in the A2A-Ising model, the stable revivals are so strong in the highly symmetric Heisenberg model considered here that the eigenvector statistics histogram plots for both 6- and 9-qubit models show clear agreement with RMT distributions in the quantum chaotic region (Figs 14 (e,f)), and clear disagreement both before the threshold (Figs 14 (b,c)) and on the stable island (Figs 14 (h,i)). By contrast, for the smallest system size in Fig. 14, none of the histogram plots provides any clear conclusion, nor are there enough eigenvector components for $X^{2}_{\rm RMT}$ to provide reliable statistical confidence.
@@ E.3 Eigenvector Statistics and the Effect of Truncated Cavity Dimension
Figure 19: Eigenvector and eigenvalue statistics of Trotterised Dicke unitary with $j=6$, and the system parameters are the same as the main text of the paper. In each panel (a–d), the main histogram plots show the eigenvector statistics, upper insets show the corresponding eigenvalue statistics, and the lower insets show the tail of the main histogram plots. Before the threshold $\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$ neither the eigenvector nor the eigenvalue statistics show an agreement with RMT distributions for cavity dimension truncations (a) $\dim_{\rm c}=25$ and (b) $\dim_{\rm c}=200$. Beyond the threshold $\tau=0.19\leavevmode\nobreak\ (2\pi g^{-1})$, both the eigenvalue and the eigenvector statistics are observed to be close to RMT distributions for (c) $\dim_{\rm c}=25$, and such agreements are lost both in the eigenvalue and the eigenvectors statists for (d) $\dim_{\rm c}=200$.
@@ E.3.1 Level-spacing and Eigenvector Statistics
Figure 19 shows the level-spacing statistics in the upper-right insets, and the eigenvector statistics in the main figures with its tail highlighted in log scale in the lower-right insets. In Figs 19 (a–b), we use the step size $\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$ (i.e. before the threshold), and as expected, we see that neither the eigenvector nor the eigenphase statistics show agreement with RMT distributions. The eigenphase statistics are actually Poisson-distributed for this step size, which is a signature of regular dynamics according to the Berry-Tabor conjecture [88]. For the step size $\tau=0.19\leavevmode\nobreak\ (2\pi g^{-1})$ (beyond the threshold) and a cavity dimension $\dim_{\rm c}=25$, both the eigenvector and eigenphase statistics show agreement with RMT distributions, shown in Fig. 19 (c). However, once $\dim_{\rm c}$ exceeds a certain value (that increases with $j$, as seen in the next section), agreement with RMT is lost in both of these cases; shown for $\dim_{\rm c}=200$ in Fig. 19 (d) where the level-spacing statistics are actually closest to a Poisson distribution. With the caveat of needing to choose an appropriate cavity truncation $\dim_{\rm c}$, eigenvector and level-spacing statistics both show corresponding agreement or disagreement with RMT beyond or before the threshold, and show the same behaviour in various different parameter regimes (not shown).
@@
Figure 20: Dependency of eigenvector statistics on cavity truncation for Trotterised Dicke models with a range of $j$ values: (a–f) showing our results are reasonably robust, and (g–i) exploring the effective RMT system dimension. (System parameters are the same as in the main text.) (a–c) $X^{2}_{\rm RMT}$ versus cavity truncation dimension $\dim_{\rm c}$ and Trotter step size $\tau$ for three different spin sizes, showing that the $X^{2}_{\rm RMT}$ threshold is roughly independent of cavity truncation dimension, with $X^{2}_{\rm RMT}$ not approaching near 1 before the threshold for any cavity truncation (shown in (d) for more spins $j$ for $\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$), and $X^{2}_{\rm RMT}\sim 1$ in a sizeable region around $\dim_{\rm c}/\dim_{j}\sim 1$ in the quantum chaotic region beyond the threshold (shown in (e) for a step size $\tau=0.19\leavevmode\nobreak\ (2\pi g^{-1})$). (f) $X^{2}_{\rm RMT}$ versus step size with the cavity truncation dimension chosen to be $\dim_{\rm c}=\lceil 0.5\times(\dim_{j}+1)\rceil$, compared with the main text of the paper, where $X^{2}_{\rm RMT}$ is presented for $\dim_{\rm c}=\dim_{j}+1$ for the same spin values. (g) Fixing the cavity truncation dimension to $\dim_{\rm c}=\dim_{j}+1=2j+2$, and comparing the resulting eigenvector statistics against COE distributions of different RMT fit dimensions $\dim_{f}$, $X^{2}_{\rm RMT}$ only reaches 1 for $\dim_{f}\sim\frac{1}{2}\dim_{j}\times\dim_{\rm c}$. (h) Comparison of the COE fit dimension $\dim_{f}$ that minimises $X^{2}_{\rm RMT}$ against the expected value of $\dim_{f}=\frac{k}{2}\dim_{j}\times\dim_{\rm c}$, and (i) the minimum value of $X^{2}_{\rm RMT}$ reached, when the cavity truncation dimension is fixed to $k\times\dim_{j}+\mod(k\dim_{j}^{2},2)$, for various $k$. (d–g,i) The shaded regions, coloured by system size with colours paler than the corresponding line colour, indicate the $99\%$ confidence regions for the chi-squared distribution, to illustrate where statistically reliable conclusions can be drawn.
@@ E.3.2 Goodness-of-fit Analyses
Next, we analyse $X^{2}_{\rm RMT}$ as a function of $\dim_{\rm c}$ for a range of $j$ values at one step size before ($\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$) and one after ($\tau=0.19\leavevmode\nobreak\ (2\pi g^{-1})$) the threshold. In this case, we determine the RMT fit dimension from the total dimension of the truncated system, i.e. $\dim_{f}=\frac{1}{2}\dim_{j}\times\dim_{\rm c}$ (accounting for parity symmetry). Figure 20 (d) shows, for a larger selection of spin sizes than above, that the eigenvector statistics before the threshold do not agree with RMT for any value of $j$ or $\dim_{\rm c}$. In the quantum chaotic region beyond the threshold, however, we observe that $X^{2}_{\rm RMT}\sim 1$ (with agreement defined by the shaded regions representing the $99\%$ confidence regions for each system size) for cavity truncations in a region around $\dim_{\rm c}/\dim_{j}$ that becomes more sharply defined with increasing $j$ (for the step size shown in Fig. 20 (e), the $X^{2}_{\rm RMT}\sim 1$ region appears to be asymptoting towards $0.3\lesssim\dim_{\rm c}/\dim_{j}\lesssim 5$), before rapidly transitioning to a region of large $X^{2}_{\rm RMT}$ for larger truncations. Though we leave it as an open question to find a rigorous explanation, our results show that there is a relatively wide range of cavity truncations that give good agreement with RMT. In Fig. 20 (f), we plot $X^{2}_{\rm RMT}$ versus step size for various system sizes, with $\dim_{\rm c}=\lceil\frac{1}{2}(\dim_{j}+1)\rceil=\lceil\frac{1}{2}(2j+2)\rceil$, showing results that are qualitatively the same as the main text, where we set $\dim_{\rm c}=\dim_{j}$ for the same spin values, with the same threshold position and overall behaviour, as further illustration that our main results regarding $X^{2}_{\rm RMT}$ for the Dicke model do not depend sensitively on $\dim_{\rm c}$.
@@ E.4 Quantum Chaos in Dicke model
The Dicke model has previously been shown to be quantum chaotic for the coupling strengths $g$ around the critical coupling $g_{c}=\sqrt{\omega_{c}\omega_{j}}/2$ of the super-radiant phase transition (it becomes regular, if $g$ is either too much smaller or larger than $g_{c}$[86, 85]. For these regimes, these papers show that the level spacing statistics of the Dicke Hamiltonian follow Wigner-Dyson statistics. They support these observations by obtaining a classical correspondence using Holstein-Primakoff transformations, together with certain approximations, and showing that the classical limit is chaotic for the parameters that give Wigner-Dyson statistics in the quantum case.

### v5

@@ C.1 Dynamics in Small Systems
Figure 12: Eigenvector statistics for Trotterised A2A-Ising unitary. Columns show increasing $j$ values from left-to-right, and the rows, from top-to-bottom, are for step sizes before the threshold ($\tau=0.002\penalty 10000\ (2\pi g^{-1})$), chaotic ($\tau=0.5\penalty 10000\ (2\pi g^{-1})$), and stable revival (seen only in large system sizes) ($\tau=0.7\penalty 10000\ (2\pi g^{-1})$) regions. The coloured lines are RMT distributions as shown by the legend.
@@ C.2 Eigenvector Statistics
In the main text of the paper, we show eigenvector statistics histograms for $j=64$ and calculate $X^{2}_{\rm RMT}$ for the other system sizes. Here, to illustrate how system size affects the histograms themselves, we provide example histograms for several $j$ values and three step sizes, which are chosen from regular (before the threshold), quantum chaotic (after the threshold), and stable revival regions. Figure 12 shows the eigenvector statistics for different $j$ values on each column and different step sizes on each row: (top to bottom) before the threshold ($\tau=0.002\penalty 10000\ (2\pi g^{-1})$), quantum chaotic ($\tau=0.5\penalty 10000\ (2\pi g^{-1})$), and stable revival (seen only in large system sizes) ($\tau=0.7\penalty 10000\ (2\pi g^{-1})$) regions, respectively. The smallest system size, in Fig. 12 (a), does not have enough components to produce meaningful statistics, and the visual comparisons in Figs 12 (b–c) do not provide very clear conclusions. The $X^{2}_{\rm RMT}$ goodness-of-fit test statistic, on the other hand, provides consistent conclusions in terms of successfully identifying the disagreement/agreement of eigenvector statistics with RMT distributions (shown in Fig. 6 (a,d)). For the large system sizes, the regular and chaotic regions are clearly distinguishable in the eigenvector statistics, shown respectively for (d) $j=64$ and (e) $j=256$ in Figs 12 (d–e). However, the stable regions ($\tau=4.5\penalty 10000\ (2\pi g^{-1})$) are visually indistinguishable from the quantum chaotic $\tau=3.5\penalty 10000\ (2\pi g^{-1})$ case, and, once again, $X^{2}_{\rm RMT}$ proves useful and identifies these weak islands (see main text).
@@ C.3 Stable Regions
The quantum-classical correspondence principle suggests the following explanation for the changes observed in the local observable in Fig. 13 (f) around $\tau=0.7\penalty 10000\ (2\pi g^{-1})$: Trotter errors are inevitably large for trajectories corresponding to chaotic dynamics that is ergodic within the entire classical phase space (as for a completely mixed state). Trajectories within the stable islands of classical phase space follow regular dynamics that is generally restricted to a more localised region, and which therefore exhibits a different overlap with the target dynamics, from that corresponding to an underlying classically chaotic trajectory. Hence the re-emergence of stable islands in the A2A-Ising’s classical phase space accompanies the revival of stability in the stable regions of the Trotterised quantum model, for initial states which overlap with the stable islands. But note again with caution that, as in the case of the Trotterised Heisenberg model, the observed revivals of stability do not typically correspond to any revival in the quality of simulation with an underlying target model.
@@ Appendix D Heisenberg Model
Figure 14: Eigenvector statistics for Trotterised Heisenberg unitary. Columns, left to right, show 3, 6, and 9 qubit chains, and rows, top-to-bottom, are for step sizes before the threshold ($\tau=2e\text{-}6\penalty 10000\ (2\pi g^{-1})$), quantum chaotic ($\tau=0.2\penalty 10000\ (2\pi g^{-1})$), and stable revival ($\tau\approx 0.5\penalty 10000\ (2\pi g^{-1})$) regions. 6 qubits and 9 qubits show clear deviations from the RMT distributions (c) before the threshold and (i) in the stable revival region, and it closely follows the COE distribution in (f) quantum chaotic region. 3 qubits cases in (a,d,g) do not have enough components to give good statistics. The coloured lines are RMT distributions as shown by the legend.
@@ D.1 Eigenvector Statistics
So far, we have only presented the processed $X^{2}_{\rm RMT}$ analysis for the Heisenberg Trotterisation eigenvector statistics. Here, Fig. 14 explicitly shows the full histogram data for 3, 6, and 9 qubits Heisenberg Trotterisations on each column and different step sizes on each row. The step sizes shown in the columns of Fig. 14 are, respectively from top-to-bottom, from before the threshold ($\tau=2e\text{-}6\penalty 10000\ (2\pi g^{-1})$), and from the quantum chaotic ($\tau=0.2\penalty 10000\ (2\pi g^{-1})$) and stable revival ($\tau\approx 0.5\penalty 10000\ (2\pi g^{-1})$) regions. (Note: Numerical diagonalisation becomes unstable at or near $\tau=0.5$ for our model parameters due to numerical imprecision, so we instead use $\tau=0.5000001$, which our analysis indicates is close enough to have “converged” to a limiting value for $\tau\rightarrow 0.5$, but far enough from $\tau=0.5$ for numerical error to be neglected.) Unlike in the A2A-Ising model, the stable revivals are so strong in the highly symmetric Heisenberg model considered here that the eigenvector statistics histogram plots for both 6- and 9-qubit models show clear agreement with RMT distributions in the quantum chaotic region (Figs 14 (e,f)), and clear disagreement both before the threshold (Figs 14 (b,c)) and on the stable island (Figs 14 (h,i)). By contrast, for the smallest system size in Fig. 14, none of the histogram plots provides any clear conclusion, nor are there enough eigenvector components for $X^{2}_{\rm RMT}$ to provide reliable statistical confidence.
@@ E.3 Eigenvector Statistics and the Effect of Truncated Cavity Dimension
Figure 19: Eigenvector and eigenvalue statistics of Trotterised Dicke unitary with $j=6$, and the system parameters are the same as the main text of the paper. In each panel (a–d), the main histogram plots show the eigenvector statistics, upper insets show the corresponding eigenvalue statistics, and the lower insets show the tail of the main histogram plots. Before the threshold $\tau=0.002\penalty 10000\ (2\pi g^{-1})$ neither the eigenvector nor the eigenvalue statistics show an agreement with RMT distributions for cavity dimension truncations (a) $\dim_{\rm c}=25$ and (b) $\dim_{\rm c}=200$. Beyond the threshold $\tau=0.19\penalty 10000\ (2\pi g^{-1})$, both the eigenvalue and the eigenvector statistics are observed to be close to RMT distributions for (c) $\dim_{\rm c}=25$, and such agreements are lost both in the eigenvalue and the eigenvectors statists for (d) $\dim_{\rm c}=200$.
@@ E.3.1 Level-spacing and Eigenvector Statistics
Figure 19 shows the level-spacing statistics in the upper-right insets, and the eigenvector statistics in the main figures with its tail highlighted in log scale in the lower-right insets. In Figs 19 (a–b), we use the step size $\tau=0.002\penalty 10000\ (2\pi g^{-1})$ (i.e. before the threshold), and as expected, we see that neither the eigenvector nor the eigenphase statistics show agreement with RMT distributions. The eigenphase statistics are actually Poisson-distributed for this step size, which is a signature of regular dynamics according to the Berry-Tabor conjecture [88]. For the step size $\tau=0.19\penalty 10000\ (2\pi g^{-1})$ (beyond the threshold) and a cavity dimension $\dim_{\rm c}=25$, both the eigenvector and eigenphase statistics show agreement with RMT distributions, shown in Fig. 19 (c). However, once $\dim_{\rm c}$ exceeds a certain value (that increases with $j$, as seen in the next section), agreement with RMT is lost in both of these cases; shown for $\dim_{\rm c}=200$ in Fig. 19 (d) where the level-spacing statistics are actually closest to a Poisson distribution. With the caveat of needing to choose an appropriate cavity truncation $\dim_{\rm c}$, eigenvector and level-spacing statistics both show corresponding agreement or disagreement with RMT beyond or before the threshold, and show the same behaviour in various different parameter regimes (not shown).
@@
Figure 20: Dependency of eigenvector statistics on cavity truncation for Trotterised Dicke models with a range of $j$ values: (a–f) showing our results are reasonably robust, and (g–i) exploring the effective RMT system dimension. (System parameters are the same as in the main text.) (a–c) $X^{2}_{\rm RMT}$ versus cavity truncation dimension $\dim_{\rm c}$ and Trotter step size $\tau$ for three different spin sizes, showing that the $X^{2}_{\rm RMT}$ threshold is roughly independent of cavity truncation dimension, with $X^{2}_{\rm RMT}$ not approaching near 1 before the threshold for any cavity truncation (shown in (d) for more spins $j$ for $\tau=0.002\penalty 10000\ (2\pi g^{-1})$), and $X^{2}_{\rm RMT}\sim 1$ in a sizeable region around $\dim_{\rm c}/\dim_{j}\sim 1$ in the quantum chaotic region beyond the threshold (shown in (e) for a step size $\tau=0.19\penalty 10000\ (2\pi g^{-1})$). (f) $X^{2}_{\rm RMT}$ versus step size with the cavity truncation dimension chosen to be $\dim_{\rm c}=\lceil 0.5\times(\dim_{j}+1)\rceil$, compared with the main text of the paper, where $X^{2}_{\rm RMT}$ is presented for $\dim_{\rm c}=\dim_{j}+1$ for the same spin values. (g) Fixing the cavity truncation dimension to $\dim_{\rm c}=\dim_{j}+1=2j+2$, and comparing the resulting eigenvector statistics against COE distributions of different RMT fit dimensions $\dim_{f}$, $X^{2}_{\rm RMT}$ only reaches 1 for $\dim_{f}\sim\frac{1}{2}\dim_{j}\times\dim_{\rm c}$. (h) Comparison of the COE fit dimension $\dim_{f}$ that minimises $X^{2}_{\rm RMT}$ against the expected value of $\dim_{f}=\frac{k}{2}\dim_{j}\times\dim_{\rm c}$, and (i) the minimum value of $X^{2}_{\rm RMT}$ reached, when the cavity truncation dimension is fixed to $k\times\dim_{j}+\mod(k\dim_{j}^{2},2)$, for various $k$. (d–g,i) The shaded regions, coloured by system size with colours paler than the corresponding line colour, indicate the $99\%$ confidence regions for the chi-squared distribution, to illustrate where statistically reliable conclusions can be drawn.
@@ E.3.2 Goodness-of-fit Analyses
Next, we analyse $X^{2}_{\rm RMT}$ as a function of $\dim_{\rm c}$ for a range of $j$ values at one step size before ($\tau=0.002\penalty 10000\ (2\pi g^{-1})$) and one after ($\tau=0.19\penalty 10000\ (2\pi g^{-1})$) the threshold. In this case, we determine the RMT fit dimension from the total dimension of the truncated system, i.e. $\dim_{f}=\frac{1}{2}\dim_{j}\times\dim_{\rm c}$ (accounting for parity symmetry). Figure 20 (d) shows, for a larger selection of spin sizes than above, that the eigenvector statistics before the threshold do not agree with RMT for any value of $j$ or $\dim_{\rm c}$. In the quantum chaotic region beyond the threshold, however, we observe that $X^{2}_{\rm RMT}\sim 1$ (with agreement defined by the shaded regions representing the $99\%$ confidence regions for each system size) for cavity truncations in a region around $\dim_{\rm c}/\dim_{j}$ that becomes more sharply defined with increasing $j$ (for the step size shown in Fig. 20 (e), the $X^{2}_{\rm RMT}\sim 1$ region appears to be asymptoting towards $0.3\lesssim\dim_{\rm c}/\dim_{j}\lesssim 5$), before rapidly transitioning to a region of large $X^{2}_{\rm RMT}$ for larger truncations. Though we leave it as an open question to find a rigorous explanation, our results show that there is a relatively wide range of cavity truncations that give good agreement with RMT. In Fig. 20 (f), we plot $X^{2}_{\rm RMT}$ versus step size for various system sizes, with $\dim_{\rm c}=\lceil\frac{1}{2}(\dim_{j}+1)\rceil=\lceil\frac{1}{2}(2j+2)\rceil$, showing results that are qualitatively the same as the main text, where we set $\dim_{\rm c}=\dim_{j}$ for the same spin values, with the same threshold position and overall behaviour, as further illustration that our main results regarding $X^{2}_{\rm RMT}$ for the Dicke model do not depend sensitively on $\dim_{\rm c}$.
@@ E.4 Quantum Chaos in Dicke model
The Dicke model has previously been shown to be quantum chaotic for the coupling strengths $g$ around the critical coupling $g_{c}=\sqrt{\omega_{c}\omega_{j}}/2$ of the super-radiant phase transition (it becomes regular, if $g$ is either too much smaller or larger than $g_{c}$) [86, 85]. For these regimes, these papers show that the level spacing statistics of the Dicke Hamiltonian follow Wigner-Dyson statistics. They support these observations by obtaining a classical correspondence using Holstein-Primakoff transformations, together with certain approximations, and showing that the classical limit is chaotic for the parameters that give Wigner-Dyson statistics in the quantum case.